THE STRUCTURE OF 1,2-DICYANOCYCLOBUTENE FROM MICROWAVE SPECTROSCOPY, ELEC-TRON DIFFRACTION AND AB INITIO CALCULATIONS.

\underline{DENIS PETITPREZ} and GEORGES WLODARCZAK, Laboratoire de Physique des Lasers, Atomes et Molecules,UMR CNRS 8523, Universite des sciences et technologies de Lille,Batiment P5,F-59655 Villeneuve d’Ascq,France; ARMIN DE MEIJERE, Institut für Organische Chemie der Georg-August-Universität Göttingen Tammannstrasse 2 D-37077 Göttingen Germany; and HARALD MOLLEN-DAL, Department of Chemistry, The University of Oslo,P. O. Box 1033 Blindern, N-0315 Oslo, Norway.

The microwave spectrum of 1,2-dicyanocyclobutene has been investigated by microwave Stark spectroscopy in the 11.0 - 34.0 GHz in Oslo. The normal species, the ^{15}N^{14}N and the ^{13}C isotopic species have been observed by molecular beam microwave Fourier transform spectroscopy in Lille. Hyperfine structures have been resolved and the quadrupole coupling constants have been determined. The molecule is found to have a symmetry plane (Cs symmetry). This structure is in good agreement with the ab initio structure, calculated at the MP2 6-311G^{∗∗} level, and the structure determined by electron diffraction experiments.